OpenStudy (anonymous):

is 2b^6 the simplified answer of (2b^2)^4

OpenStudy (campbell_st):

no... your problem uses the power of a power rule \[(x^a)^b = x^{a \times b}\] so you would have \[(2b^2)^4 = 2^{(1 \times 4)}b^{(2 \times 4)}\] hope it helps

OpenStudy (anonymous):

ohh i see now thank you it did help